# Robust parameter estimation for partially observed second-order diffusion processes

### v1

@@ Abstract.
Estimating parameters of a diffusion process given continuous-time observations of the process via maximum likelihood approaches or, online, via stochastic gradient descent or Kalman filter formulations constitutes a well-established research area. It has also been established previously that these techniques are, in general, not robust to perturbations in the data in the form of temporal correlations. While the subject is relatively well understood and appropriate modifications have been suggested in the context of multi-scale diffusion processes and their reduced model equations, we consider here an alternative setting where a second-order diffusion process in positions and velocities is only observed via its positions. In this note, we propose a simple modification to standard stochastic gradient descent and Kalman filter formulations, which eliminates the arising systematic estimation biases. The modification can be extended to standard maximum likelihood approaches and avoids computation of previously proposed correction terms.
@@ 1. Introduction
from incoming observations $X_{t}^{\dagger}\in\mathbb{R}^{d}$, $t\geq 0$, of an underlying reference process. Here $f:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ and $F:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}^{d\times d_{\theta}}$ are given functions, $\sigma>0$ is the diffusion constant, and $W_{t}$ denotes $d$-dimensional Brownian motion.
The data $X_{t}^{\dagger}$ is being generated from a second-order SDE of the same type, i.e.,
@@
with the drift $g(x,u)$ being unknown. We assume that there exists a unique $\theta_{\ast}$ such that
@@
While standard online estimation techniques, such as stochastic gradient descent (SGD) [17] and Kalman filtering [11, 16], can be applied in case both $(X_{t}^{\dagger},U_{t}^{\dagger})$ from (2) are being observed, the estimation problem becomes more challenging in case only positions $X_{t}^{\dagger}$ are being observed and velocities need to be estimated using finite differencing [15, 5], i.e.,
@@
As demonstrated previously [9], naive application of standard maximum likelihood-based (MLE) estimators [10, 13] leads to a systematic bias. In order to correct for the arising estimation bias, appropriate corrections terms to the MLE estimator have been proposed in [9, 5, 3]. Related phenomena have been investigated in the context of multi-scale diffusion [2, 12, 4], which again can be corrected via an appropriately modified MLE estimator [8].
@@
As demonstrated in the next section by means of a simple example, a discretised SGD of the form
@@
denotes the learning rate for appropriate constants $c_{i}>0$, $i=1,2$.
@@ 2. Motivating example
(10) $X_{t}^{\dagger}=X_{s}^{\dagger}+(t-s)U_{s}^{\dagger}+\sqrt{\sigma}\int_{s}^{t}W_{r-s}{\rm d}r.$
Here we have introduced the shorthand $W_{t-s}^{\dagger}=W_{t}^{\dagger}-W_{s}^{\dagger}$.
@@
(12) $I_{t}=\int_{0}^{t}\left({\rm d}U_{s}^{\dagger}+\theta_{s}U_{s}^{\dagger}{\rm d}s\right)=\int_{0}^{t}\left((\theta_{s}-\theta_{\ast})U_{s}^{\dagger}{\rm d}t+\sqrt{\sigma}{\rm d}W_{s}^{\dagger}\right)$
@@
In terms of numerical implementations of the SGD formulation 11), we consider the discrete time SGD method
@@
Using (6) instead of the true $U_{t_{n}}$ etc. lead to the previously stated (5), which reduces here to
@@
This violation of the martingale property is perhaps not surprising since (19) resembles a discretised Stratonovitch integral [13]. We note that such an interpretation as a discretised stochastic integral concerns only the second-order differencing term in (19) and we may therefore set $\theta=\theta_{\ast}=0$ in the subsequent derivations. Recall that
@@
(21a) $\frac{1}{2}\mathbb{E}\left[\left(\tilde{U}_{n+1/2}+\tilde{U}_{n-1/2}\right)(\tilde{U}_{n+1/2}-\tilde{U}_{n-1/2})\right]$ $=\frac{1}{2}\mathbb{E}\left[\tilde{U}_{n+1/2}^{2}-\tilde{U}_{n-1/2}^{2}\right]$
@@
Hence let us analyse the statistical properties of $\tilde{\Xi}_{n}$. We find that
@@
(25) $\mbox{var}\,\left(\int_{t_{n}}^{t_{n+1}}W_{t-t_{n}}^{\dagger}{\rm d}t\right)=\frac{1}{3}\tau^{3}.$
@@
In line with the previous result [9], our calculations imply the following estimator for the diffusion constant $\sigma$:
@@
which implies that $\sqrt{\tau}\tilde{\Xi}_{n}$ converges to standard Brownian motion in the limit $\tau\to 0$. When using this result in the SGD formulation (22) for general $\theta_{\ast}$, we obtain
@@
3. Bias-corrected SGD
The analysis from Section 2 leads us to the following modification of the discrete-time SGD implementation (5):
@@ 3. Bias-corrected SGD
We note that unbiased formulations are not unique. For example, (32) could be replaced by
@@
This freedom is also utilised when formulating the modified MLE estimator (37) further below in Section 6.
4. Bias-corrected Kalman filter
The SGD method (32) provides an online point estimator for the parameter $\theta$. However, it is also possible to implement an online Bayesian estimator using the Kalman filter [11, 16]. The key idea is to treat the unknown parameter as a Gaussian random variable $\Theta_{n}$ with mean $m_{n}$ and covariance $\Sigma_{n}$, $n\geq 0$. The evolution equation for the mean and the covariance follow the standard Kalman update formulas:
@@ 4. Bias-corrected Kalman filter
for $n\geq 1$. The iteration is initialised from the Gaussian prior $\Theta_{1}\sim{\rm N}(m_{\rm prior},\Sigma_{\rm prior})$.
In contrast to the SGD formulation (32), the learning rate $\alpha_{t_{n}}$ no longer appears and has instead been submerged implicitly into the Kalman gain $K_{n}$. Indeed, the singular values of the covariance matrix $\Sigma_{n}$ decay with rate $1/t_{n}$ under suitable ergodicity assumptions on (2). If the required diffusion constant $\sigma$ is unknown, it can either be estimated via (28) or, alternatively, be treated as part of the prior covariance matrix $\Sigma_{\rm prior}$ and then set to $\sigma=1$ in the Kalman gain (35).
@@ 5. Numerical example
We consider the nonlinear second-order SDEs
@@
with unknown parameter $\theta\in\mathbb{R}$. The data $X_{t}^{\dagger}$ is generated with $\theta_{\ast}=1$, while the SGD iteration starts from $\theta_{0}=2$. The sample interval is set to $\tau=0.025$. The data is generated with a much smaller step-size of $h=\tau/100$. A total of $N=10^{5}$ positions $X_{t_{n}}$, $n=1,\ldots,N$ have been generated. The learning rate is given by $\alpha_{t_{n}}=0.2/n$. See Figure 1 for the time evolved estimates $\theta_{n}$, which clearly converge to $\theta_{\ast}$. We also display results from the standard SGD in Figure 1. While the standard SGD converges slightly faster, it converges to a wrong parameter value.
Figure 1. Left panel: Estimated parameter value from standard SGD as a function of the iteration index. The iteration starts at $\theta_{0}=2$ and the reference value is $\theta_{\ast}=1$. A systematic bias can clearly be seen. Right panel: Same data is used in the modified SGD. The modified SGD algorithm converges to the correct reference value.
@@ 6. Conclusions
Future directions include the investigation of the proposed estimators in the context multi-scale diffusion when (1) arises from homogenisation of an underlying multi-scale diffusion process and the inclusion of observation errors in the positions $X_{t}$. In the former case, a combination of the approaches taken here and in [14] emerges as a feasible starting point.
Acknowledgements.

### v2

# Parameter estimation for partially observed second-order diffusion processes

@@ Abstract.
Estimating parameters of a diffusion process given continuous-time observations of the process via maximum likelihood approaches or, online, via stochastic gradient descent or Kalman filter formulations constitutes a well-established research area. It has also been established previously that these techniques are, in general, not robust to perturbations in the data in the form of temporal correlations of the driving noise. While the subject is relatively well understood and appropriate modifications have been suggested in the context of multi-scale diffusion processes and their reduced model equations, we consider here an alternative but related setting where a diffusion process in positions and velocities is only observed via its positions. In this note, we propose a simple modification to standard stochastic gradient descent and Kalman filter formulations, which eliminates the arising systematic estimation biases. The modification can be extended to standard maximum likelihood approaches and avoids computation of previously proposed correction terms.
@@ 1. Introduction
from incoming observations $X_{t}^{\dagger}\in\mathbb{R}^{d}$, $t\geq 0$, of an underlying reference process. Here $f:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ and $F:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}^{d\times d_{\theta}}$ are given functions, $\sigma>0$ is the diffusion constant, and $W_{t}$ denotes $d$-dimensional Brownian motion. The data $X_{t}^{\dagger}$ is being generated from a second-order SDE of the same type, i.e.,
@@
with the drift $g(x,u)$ being unknown.
Note that for $f,F,g$ independent of position, the system corresponds to a first-order diffusion process where only the integral of the process is being observed.
We assume that $g(x,u)$ is such that the reference process $(U_{t}^{\dagger},X_{t}^{\dagger})$ is ergodic with invariant distribution $\pi$. The optimal parameter $\theta_{\ast}$ can then be determined by minimizing the ergodic least squares functional
(3) $\mathcal{L}(\theta)=\frac{1}{2}\mathbb{E}_{\pi}\left[\|g(X,U)-f(X,U)-F(X,U)\theta\|^{2}\right],$
which has a unique minimizer provided the Fisher information matrix $\mathbb{E}_{\pi}[F(X,U)^{T}F(X,U)]$ has full rank. Moreover, we assume that the unique minimizer satisfies
@@
However, the gradient of $\mathcal{L}(\theta)$ cannot be computed since $g(x,u)$ is unknown. Therefore we investigate online estimation techniques in this paper based on observed positions $X_{t}^{\dagger}$.
While standard online estimation techniques, such as stochastic gradient descent (SGD) [21] and Kalman filtering [14, 19], can be applied in case both ($X_{t}^{\dagger},U_{t}^{\dagger})$ from (2) are being observed, the estimation problem becomes more challenging in case only positions $X_{t}^{\dagger}$ are being observed and velocities need to be estimated using finite differencing [18, 6], such as,
@@
As demonstrated previously [10, 12, 11], naive application of standard maximum likelihood-based (MLE) estimators [13, 16] leads to a systematic bias. In order to correct for the arising estimation bias, appropriate corrections terms to the MLE estimator have been proposed in [12, 6, 3, 11]. An alternative approach has been to treat the continuous time velocity process as a latent variable and calculate the MLE estimator using an expectation-maximization algorithm [4]. Viewing the SDEs as a partially observed system, the likelihood of $\theta$ could also be approximated using an extended Kalman filter [20]. Related phenomena have been investigated in the context of multi-scale diffusion [2, 15, 5], which again can be corrected via an appropriately modified MLE estimator [9].
@@
As demonstrated in the next section by means of a simple example, a discretized SGD of the form
@@
denotes the learning rate for appropriate constants $c_{i}>0$, $i=1,2$. While the centered approximation in (LABEL:eq:U_symmetric_is_the_best_approximation_to_\$U_n\$_from_the_position_measurements_at_the_nearest_neighbor_time_points_\$X_t_n-1,_X_t_n,X_t_n+1\$_alternative_approximations_to_the_velocity_at_\$t_n\$_could_also_be_employed.)
@@ 2. Motivating example
(11) $X_{t}^{\dagger}=X_{s}^{\dagger}+(t-s)U_{s}^{\dagger}+\sqrt{\sigma}\int_{s}^{t}W^{\dagger}_{r-s}{\rm d}r.$
Here, we have introduced the shorthand $W_{t-s}^{\dagger}=W_{t}^{\dagger}-W_{s}^{\dagger}$.
@@
(13) $I_{t}=\int_{0}^{t}\left({\rm d}U_{s}^{\dagger}+\theta_{s}U_{s}^{\dagger}{\rm d}s\right)=\int_{0}^{t}\left((\theta_{s}-\theta_{\ast})U_{s}^{\dagger}{\rm d}s+\sqrt{\sigma}{\rm d}W_{s}^{\dagger}\right)$
@@
In terms of numerical implementations of the SGD formulation (12), we consider the discrete time SGD method
@@
Using (7) instead of the true $U_{t_{n}}^{\dagger}$ etc.  and symmetrizing the gain leads to the previously stated (6), which reduces here to
@@
This violation of the martingale property is perhaps not surprising since (20) does not resemble a discretized Itô-Integral like (16) but rather a discretized Stratonovitch integral [16]. We note that such an interpretation as a discretized stochastic integral concerns only the second-order differencing term in (20) and we may therefore set $\theta=\theta_{\ast}=0$ in the subsequent derivations. Recall that
@@
(22a) $\frac{1}{2}\mathbb{E}\left[\left(\tilde{U}_{n+1/2}+\tilde{U}_{n-1/2}\right)\left(\tilde{U}_{n+1/2}-\tilde{U}_{n-1/2}\right)\right]$ $=\frac{1}{2}\mathbb{E}\left[\tilde{U}_{n+1/2}^{2}-\tilde{U}_{n-1/2}^{2}\right]$
@@
Note that the innovation in the modified implementation (23) only contains positional data from time points that are larger than or equal to those in the gain.
@@
Hence let us analyze the statistical properties of $\tilde{\Xi}_{n}$. We find that
@@
(26) $\mbox{var}\,\left(\int_{t_{n}}^{t_{n+1}}W_{t-t_{n}}^{\dagger}{\rm d}t\right)=\frac{1}{3}\tau^{3}$
@@
which implies that
(30) $Z_{t}\mathrel{\mathop{:}}=\begin{cases}0&\text{if}\quad t<\tau\\ \sum_{k=1}^{\lfloor t/\tau\rfloor}\sqrt{\tau}\,\tilde{\Xi}_{k}&\text{else}\end{cases}$
converges to standard Brownian motion in the limit $\tau\to 0$. When using this result in the SGD formulation (23) for general $\theta_{\ast}$, we obtain
@@
Remark: Independent of the drift estimation, it is well known that the diffusion term can be estimated from the variability of a process. In line with previous results [7, 11, 12], our above calculations and equation (28) imply the following estimator for the diffusion constant $\sigma$ from the quadratic variation of the approximated velocities $\tilde{U}_{n+1/2}$:
3. Unbiased SGD
The analysis from Section 2 leads to the following modification of the discrete-time SGD implementation (6):
@@ 3. Unbiased SGD
We note that unbiased formulations are not unique since only the shifting of the second-order differecing term in the innovation is relevant for consistency. For example, (34) could be replaced by
@@
This freedom is also utilized when formulating the modified MLE estimator (39) further below in Section 6.
4. Unbiased Kalman filter
The SGD method (34) provides an online point estimator for the parameter $\theta$. However, it is also possible to implement an online Bayesian estimator using the Kalman filter [14, 19]. The key idea is to treat the unknown parameter as a Gaussian random variable $\Theta_{n}$ with mean $m_{n}$ and covariance matrix $\Sigma_{n}$ and with trivial dynamics $\Theta_{n+1}=\Theta_{n}$, $n\geq 0$ [20]. The evolution equation for the mean and the covariance matrix follow the standard Kalman update formulas:
@@ 4. Unbiased Kalman filter
for $n\geq 1$. The iteration is initialized from the Gaussian prior $\Theta_{1}\sim{\rm N}(m_{\rm prior},\Sigma_{\rm prior})$.
In contrast to the SGD formulation (34), the learning rate $\alpha_{t_{n}}$ no longer appears and has instead been submerged implicitly into the Kalman gain $K_{n}$. Indeed, the singular values of the covariance matrix $\Sigma_{n}$ decay with rate $1/t_{n}$ under the already stated ergodicity assumption on (2) and the non-singularity of the Fisher information matrix $\mathbb{E}_{\pi}\left[F(X,U)^{\rm T}F(X,U)\right]$. If the required diffusion constant $\sigma$ is unknown, it can either be estimated via (33) or, alternatively, be treated as part of the prior covariance matrix $\Sigma_{\rm prior}$ and then set to $\sigma=1$ in the Kalman gain (37).
@@ 5. Numerical example
We consider the nonlinear second-order SDE
@@
with unknown parameter $\theta\in\mathbb{R}$. The data $X_{t}^{\dagger}$ is generated with $\theta_{\ast}=1$, while the SGD iteration starts from $\theta_{0}=2$. The sample interval is set to $\tau=0.025$. The data is generated with a much smaller step-size of $h=\tau/100$. A total of $N=10^{5}$ positions $X_{t_{n}}$, $n=1,\ldots,N$ have been generated. The learning rate is given by $\alpha_{t_{n}}=6.0/n$. See Figure 1 middle panel for the time evolved estimates $\theta_{n}$ generated by the modified SGD (34). These estimates clearly converge to $\theta_{\ast}$. In Figure 1 we also display results from the standard SGD as well as the modified Kalman filter presented in section 4. For the Kalman filter, a Gaussian prior was used with $m_{\rm{prior}}=2$ and $\Sigma_{\rm{prior}}=6$. The diffusion constant $\sigma$ was assumed to be unknown and treated as part of the prior. While the standard SGD converges slightly faster than the modified SGD, it converges to a wrong parameter value. The estimates of the modified Kalman filter are very similar to estimates of the modified SGD. An unmodified Kalman filter (not shown) on the other hand leads to a similar estimation bias as the standard SGD.
Figure 1. Left panel: Estimated parameter value from standard SGD as a function of the iteration index. The iteration starts at $\theta_{0}=2$ and the reference value is $\theta_{\ast}=1$. A systematic bias can clearly be seen. Middle panel: Same data is used in the modified SGD. The modified SGD algorithm converges to the correct reference value. Right panel: Same data is used in the unbiased Kalman filter with prior distribution $N(2,1)$ and $\sigma$ treated as part of the prior. The lightly colored lines indicate the $1\sigma$-intervals of the Bayesian estimates.
@@ 6. Conclusions
Compare this likelihood function to the contrast function presented in [12], which is based on adding an appropriate correction term. Please also note that (39a)-(39b) replaces the ergodic least squares functional (3) using measured positions and estimated velocities and that (39c) delivers the estimator (33).
Future directions include the investigation of the proposed estimators in the context multi-scale diffusion when (1) arises from homogenization of an underlying multi-scale diffusion process and the inclusion of observation errors in the positions $X_{t}$. In the former case, a combination of the approaches taken here and in [17] emerges as a feasible starting point.
Acknowledgments.